# Deep Instruction Tuning for Segment Anything Model

### v1

@@ 4.4.2 Ablation Study
The effect of freezing encoders. In Tab. LABEL:backbone, we also ablate DITs under the parameter efficient setting (Zhou et al., 2022a, b; Luo et al., 2023). It can be seen that when freezing the SAM’s image backbone and the directly plugged-in BERT, L-DIT can still achieve an oIoU of 67.16 on RefCOCO val, which is even much better than fully tuning the default SAM, as shown in Tab. LABEL:Ablation_of_Tuning_method. Unfreezing both encoders can greatly improve performance. But compared with the pre-trained BERT, the significance of updating image encoder is much more obvious, e.g., +4.88 oIoU on val. This result suggests that in DIT, the image encoder is capable of visual feature learning and cross-modal interactions, well confirming our intuition about DIT.
Figure 4: Predictions of the default SAM and our end-to-end and layer-wise DITs, i.e., E-DIT and L-DIT. Both methods can follow the text instruction better than the default SAM, while L-DIT can accurately segment the referent even with complex text content or image background. For better comparison, we use red boxes to ground the failed segmentations.
@@
Different text word injections. In our layer-wise DIT, we project the text prompts directly onto the visual space of each SAM’s layer. Here, we also test DIT with alternative manners in Tab. LABEL:project, such as cross-attention and cross-attention+FFN. Interestingly, these complex methods do not further enhance SAM’s text-aware capability, but declining performance to some extent. This case suggests that with the strong dependency modeling of SAM’s ViT encoder, a direct semantic projection can well facilitate cross-modal alignment, well confirming the motivation of our DIT.
4.5 Comparison with the State-of-the-art
@@
In Tab. 5, we compare our L-DIT against the state-of-the-art (SOTA) methods of referring image segmentation on RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016) datasets using the oIoU metric. Specially, on RefCOCO dataset, we outperforms existing trained from scratch SOTA methods with a absolute improvements of 1.94%, 1.56%, 2.91% in all three splits and performs on par with pretrained PolyFormer (Liu et al., 2023d). Furthermore, on RefCOCO+ and RefCOCOg datasets, the L-DITH also exhibits exceptional performance, demonstrating comparable capabilities to SOTAs (Liu et al., 2023d; Yang et al., 2023). In addition, compared with SOTA RIS methods using a large number of visual grounding data for pre-training, Our L-DITH is also very competitive and even achieves new SOTA performance on RefCOCO val.
4.6 Qualitative Analysis
@@
To gain deep insights into DIT, we visualize the predictions of our DIT tuning methods and SAM in Fig.4. In the first column, it becomes evident that the default SAM struggles with a nuanced understanding of text, often yielding predictions that inadequately encapsulate the referent or outright incorrect. In contrast, the simple E-DIT offers a notable improvement in SAM’s text comprehension capabilities. For instance, in the third example, SAM’s segmentation exhibit a pronounced bias towards the distracting elements on the left, while E-DIT rectifies this issue, resulting in more accurate prediction. Notably, through E-DIT enhances the performance in some cases, it can lead to the loss of certain semantic information embedded within the textual features. As a consequence, E-DIT can occasionally misinterpret the segmentation target, as illustrated in the last example. The last column reveals that L-DIT effectively mitigates these issues without necessitating specific designs for cross-modal features. Remarkably, even without providing explicit image positions, akin to the original SAM, L-DIT consistently delivers superior segmentation results. This demonstrates that L-DIT processing can effectively harness positional information from both modalities and maintain a strong degree of context-awareness, improving accuracy and robustness.
5 Conclusion
@@
In this paper, we propose two deep instruction tuning (DIT) methods to address the weakness of SAM on following text instructions. One is end-to-end and the other is layer-wise. DITs regard SAM’s default visual encoder as a stand-alone multi-modal learner instead of adding additional fusion branches. Both approaches contribute to enhancing SAM’s ability to position segmented objects based on textual guidance. To validate DITs, we conduct extensive experiments on three RIS benchmark datasets. Experimental results show that our simple DIT methods can greatly improve the capabilities of SAM. Meanwhile, the proposed layer-wise DIT can even help SAM compete with a set of SOTA methods of RIS on all three datasets.
6 Broader Impact
This paper reveals the key drawbacks of SAM in text-guided segmentation, and propose two simple yet effective deep instruction tuning methods for SAM. The findings and methods of this paper can further advance the development of SAM in the machine learning community. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ 6 Broader Impact
Acknowledgements. This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).

### v2

@@ 4.3.4. The effect of freezing encoders
In Tab. LABEL:backbone, we also ablate DITs under the parameter efficient setting (Luo et al., 2023; Wu et al., 2023e). It can be seen that when freezing the SAM’s image backbone and the directly plugged-in BERT, L-DIT can still achieve 67.16% oIoU on RefCOCO val, which is even much better than fully tuning the default SAM, as shown in Tab. LABEL:Ablation_of_Tuning_method. Unfreezing both encoders can greatly improve performance. But compared with the pre-trained BERT, the significance of updating image encoder is more obvious, e.g., +4.88% oIoU on val. This result suggests that in DIT, the image encoder is capable of visual feature learning and cross-modal interactions, well confirming our intuition about DIT.
4.3.5. Different text word injections
@@
In our layer-wise DIT, we project the text prompts directly onto the visual space of each SAM’s layer. Here, we also test DIT with alternative manners in Tab. LABEL:project, such as cross-attention and cross-attention+FFN. Interestingly, these complex methods do not further enhance SAM’s text-aware capability, but decline performance to some extent. This case suggests that with the strong dependency modeling of SAM’s ViT encoder, a direct semantic projection can already facilitate cross-modal alignment, well confirming the motivation of our DIT.
Figure 4. Predictions of the default SAM and our end-to-end and layer-wise DITs, i.e., E-DIT and L-DIT. Both methods can follow text instructions better than the default SAM, while L-DIT can accurately segment the referent even with complex text content or image background. For better comparison, we use red boxes to ground the incorrect segmentations.
4.3.6. The impact of tuning on SAM
In Tab. LABEL:tuning, we further examine the default capability of SAM with point and box guidances on the RefCOCO benchmark. Here, we use the ground-truth boxes of RefCOCO for box-guided SAM, and randomly sample 5 points on the ground-truth mask for point-guided segmentation. From Tab. LABEL:tuning, we can first see that on RefCOCO, the default SAM is still likely to segment the incorrect regions of the referent with ground-truth boxes. To explain, RefCOCO is a challenging benchmark, of which images often involve dense visual semantics, and a large box will make SAM prone segment irrelevant region. In contrast, the point sampled from the object offers a better guidance, but it still lags behind SOTA performance of RIS (Liu et al., 2023c, f). Meanwhile, the joint tuning with all types of prompts can further improve the performance of SAM on following boxes and points. We also notice that our DIT method can help SAM achieve strong text-guided performance that is already close to the box-guided capability, well validating the effectiveness of our method. In terms of the inference speed, since the number of text tokens is often about 15-20, much fewer than that of image patches, the overall impact of DITs on inference time is limited.
4.3.7. Comparison with the State-of-the-art
@@
In Tab. 7, we compare our L-DIT against the state-of-the-art (SOTA) methods of referring image segmentation on RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016) datasets using the oIoU metric. Specially, on RefCOCO dataset, we outperforms existing SOTA methods trained from scratch with absolute improvements of 1.94%, 1.56%, 2.91% in all three splits and performs on par with the pretrained PolyFormer (Liu et al., 2023c). Furthermore, on RefCOCO+ and RefCOCOg datasets, the L-DITH also exhibits exceptional performance, demonstrating comparable capabilities to SOTAs (Liu et al., 2023c; Yang et al., 2023). In addition, compared with SOTA RIS methods using a large number of visual grounding data for pre-training, Our L-DITH is very competitive and achieves new SOTA performance on RefCOCO val.
4.4. Qualitative Analysis
@@
To gain deep insights into DIT, we visualize the predictions of our DIT tuning methods and the default SAM in Fig.4. In the first column, it becomes evident that the default SAM struggles with a nuanced understanding of text, often yielding predictions that inadequately encapsulate the referent or outright incorrect. In contrast, the simple E-DIT offers a notable improvement of SAM’s text comprehension capabilities. For instance, in the third example, SAM’s segmentation exhibits a pronounced bias towards the distracting elements on the left, while E-DIT rectifies this issue, resulting in more accurate prediction. Notably, through E-DIT enhances the performance in some cases, it can lead to the loss of certain semantic information embedded within the textual features. As a consequence, E-DIT occasionally misinterprets the segmentation target, as illustrated in the last example. The last column reveals that L-DIT effectively mitigates these issues without necessitating specific designs for cross-modal interactions. Remarkably, even without providing explicit image positions, akin to the original SAM, L-DIT consistently delivers superior segmentation results. This demonstrates that L-DIT can effectively harness positional information from both modalities and maintain a strong degree of context-awareness, improving accuracy and robustness.
5. Conclusion
@@
In this paper, we propose two deep instruction tuning (DIT) methods to address the weakness of SAM on following text instructions, one is end-to-end and the other is layer-wise. DITs regard SAM’s default visual encoder as a stand-alone multi-modal learner instead of adding additional fusion branches. Both approaches contribute to enhancing SAM’s ability of segmenting target objects based on textual guidance. To validate DITs, we conduct extensive experiments on three RIS benchmark datasets. Experimental results show that our simple DIT methods can greatly improve the capability of SAM. Meanwhile, the proposed layer-wise DIT can even help SAM compete with a set of SOTA methods of RIS on all three datasets.
Acknowledgements.
@@
This work was supported by National Science and Technology Major Project (No. 2022ZD0118201), the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No.2021J01002, No.2022J06001).
Appendix A Quantitative Analysis
Figure 5. The attention maps of the default SAM and our end-to-end and layer-wise DITs, i.e., E-DIT and L-DIT. Both methods can follow text instructions better than the default SAM and adjust the area of interest according to text instructions.
A.1. Training Memory Consumption
We compare the memory overhead of L-DIT and existing methods in Tab.  LABEL:memory. From this table, we can see that the additional cost of backbone frozen is close to linear probe tuning and smaller than the full finetuning. In addition, the GPU overhead of full finetuning is even much smaller than LAVT (Yang et al., 2022) in referring image segmentation, since L-DIT does not require another deep fusion model. Compared with LAVT, our L-DIT is simpler and more intuitive, which only requires linear projections to insert text words into each SAM’s layer. These results further validate the efficiency of L-DIT.
Table 8. The memory consumption of L-DIT and existing methods. Here, ✓means that the encoder is frozen, while ✗is updated.
Model Visual Text Batch Size GPU memory
L-DIT 32 11.9G
L-DIT 32 12.4G
L-DIT 32 17.1G
LAVT (Yang et al., 2022) 8 24G
ReLA (Liu et al., 2023a) 32 12G
A.2. The impact of shared parameters for the injection of text features into the visual backbone
In the visual backbone, the shallow and deep layers typically extract low- and high-level features, respectively. To validate this intuition, we experiment on the progressive cascade structure in Tab. LABEL:cascade. From this table, we see that there is no obvious difference between layer-wise and cascade structures. To explain, DIT aims to project the text words onto the visual semantic space of SAM, and simple linear projections may be enough. Similar findings can be also found on recent MLLMs (liu2024visual), where the bridge part is often a stack of linear layers.
Table 9. The impact of shared parameters for the injection of text features into the visual backbone. Cascade means linear layers replace with progressive cascade structure.
Structure val testA testB
mIoU oIoU mIoU oIoU mIoU oIoU
Layer-wise DIT 69.97 67.50 72.44 69.55 68.12 64.77
Cascade DIT 69.94 66.76 72.20 69.72 66.93 63.38
Appendix B Qualitative Analysis
To gain deep insights into DIT, we visualize the attention maps of our DIT tuning methods and the default SAM in Fig. 5. We observe some interesting findings about the self-attention of SAM backbone after inserting text tokens.
For instance, under the frozen backbone setting, the inserted text tokens mainly interact with the visual ones on the higher layers of SAM’s encoder, e.g., the 8-th or 11-th layers. This suggests that the text mainly interacts with the high-level visual semantics of SAM when only tuning the DIT projections.
In contrast, under the full tuning setting, the interactions mainly happen at the shallow layers, e.g., 2-nd and 4-th layers. It might indicate that the text information is early embedded into the visual tokens and then for the deep multi-modal fusion in the image encoder of SAM.
Moreover, it can be observed from the visual attention map that for each layer of the visual backbone, the attentive area of the text instruction is different. For default SAM, it has no obvious interested area, which mainly divides the image into foreground and background. However, for E-DIT and L-DIT, it can be clearly seen that each layer has its attention tendency for different text instructions. The initial layers are interested in entity nouns in text instructions. The 4-th layer is more sensitive to comparative words. At 11-th layer, it mainly focuses on digital information. The 12-th layer is more inclined to color information.